# Curvature pinching of asymptotically conical gradient expanding Ricci solitons

### v2

@@ 5.3. Additional curvature pinching results
    Suppose $(M^{n},g,f)$ has WPIC1 and that the asymptotic cone has positive scalar curvature and WPIC2. Then, $(M^{n},g,f)$ must have WPIC2.
@@
    Suppose $(M^{n},g,f)$ is Kähler and has nonnegative orthogonal bisectional curvature. If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
@@
(a) By the same computations as in [72, Theorem 27] (see also [38, Lemma 2.4]), the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the Riemann curvature operator is satisfied.
@@
(c) By the same computations as in [70, Proposition 4.2], the differential inequality (5.1) in Lemma 5.1 for the sum of the two least eigenvalues of the Ricci tensor is satisfied.
(d) First of all, by the same computations as in [70, Lemma 6.1], the differential inequality (5.1) in Lemma 5.1 for the least bisectional curvature is satisfied. Hence, by Lemma 5.1, the expanding Kähler-Ricci soliton has nonnegative bisectional curvature. Moreover, by the same arguments as in [71, Theorem 3.3], one can show that the smallest eigenvalue of the complex sectional curvature also satisfies the differential inequality (5.1) in Lemma 5.1.
(e) Finally, by the same computations as in [43, Theorem 3.2] for $n\geq 12$ and in [40] for $9\leq n\leq 11$, the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied. ∎
@@
Other results that have been derived using B.-L. Chen’s lemma (see, e.g., [43, Corollary 2.4] or [25, Lemma 2.6]) for ancient solutions can similarly be extended to the setting of asymptotically conical gradient expanding Ricci solitons, following the same approach as above.
6. Appendix
@@
$W_{1a1b}$ $=-\frac{1}{(n-2)r^{2}}\left(\bar{R}_{ab}-\frac{\bar{R}}{(n-1)}\bar{g}_{ab}\right),\quad 2\leq a,b\leq n,$ (6.4)
@@
6.1. Curvature tensor of cones with vanishing Weyl tensor
@@
where $(\Sigma^{n-1},\bar{g})$ is a closed (n1)(n-1)-dimensional Riemannian manifold. Then, by (6.4), we see that $\mathcal{C}$ has vanishing Weyl curvature $W=0$ if and only if its link $\Sigma$ is a space form. Thus, we immediately have the following
Lemma 6.2.
@@
6.2. Curvature decomposition and curvature operator of 4D cones
@@
$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$, (6.5)
@@
$\langle X\wedge Y,V\wedge W\rangle=\langle X,V\rangle\langle Y,W\rangle-\langle X,W\rangle\langle Y,V\rangle.$
Observe that, for the matrices $A$ and $C$ in (6.5), we have
@@
so that $\{e_{1},e_{2},e_{3},e_{4}\}$ forms an orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, by (6.1), we have
@@
Therefore, by (6.6), we obtain
$A_{11}=C_{11}$ $=\tfrac{1}{2}R_{3434}=\tfrac{1}{2r^{2}}(m_{3}-1),$ (6.8)
@@
Similarly, by (6.7), we have
$B_{11}$ $=-\tfrac{1}{2}R_{3434}=-\tfrac{1}{2r^{2}}(m_{3}-1)=-A_{11},$ (6.9)
@@
$R_{11}$ $=0,$ $R_{22}=r^{-2}\left(m_{1}+m_{2}-2\right),$
$R_{33}$ $=r^{-2}\left(m_{1}+m_{3}-2\right),$ $R_{44}=r^{-2}\left(m_{2}+m_{3}-2\right).$
@@
    $(\mathcal{C}^{4},g_{c})$ has half WPIC if and only if it has WPIC and nonnegative Ricci curvature $Rc\geq 0$, if and only if $\overline{Rc}\geq 2{\bar{g}}$.

### v3

@@ 5.3. Additional curvature pinching results
    Suppose $(M^{n},g,f)$ has WPIC1555See, e.g., [3, Page 97] or [6, 11] for the definitions of WPIC1 and WPIC2. and that the asymptotic cone has positive scalar curvature and WPIC2. Then, $(M^{n},g,f)$ must have WPIC2.
@@
    Suppose $(M^{n},g,f)$ is Kähler and has nonnegative orthogonal bisectional curvature [^6] See, e.g., [71, Page 28] for the definition of nonnegative orthogonal bisectional curvature.. If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
@@
(a) By the same computations as in [73, Theorem 27] (see also [40, Lemma 2.4]), the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the Riemann curvature operator is satisfied.
@@
(c) By the same computations as in [71, Proposition 4.2], the differential inequality (5.1) in Lemma 5.1 for the sum of the two least eigenvalues of the Ricci tensor is satisfied.
(d) First of all, by the same computations as in [71, Lemma 6.1], the differential inequality (5.1) in Lemma 5.1 for the least bisectional curvature is satisfied. Hence, by Lemma 5.1, the expanding Kähler-Ricci soliton has nonnegative bisectional curvature. Moreover, by the same arguments as in [72, Theorem 3.3], one can show that the smallest eigenvalue of the complex sectional curvature also satisfies the differential inequality (5.1) in Lemma 5.1.
(e) Finally, by the same computations as in [44, Theorem 3.2] for $n\geq 12$ and in [42] for $9\leq n\leq 11$, the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied. ∎
@@
Other results that have been derived using B.-L. Chen’s lemma (see, e.g., [44, Corollary 2.4] or [27, Lemma 2.6]) for ancient solutions can similarly be extended to the setting of asymptotically conical gradient expanding Ricci solitons, following the same approach as above.
Appendix A Curvature properties of cones
@@
$W_{1a1b}$ $=-\frac{1}{(n-2)r^{2}}\left(\bar{R}_{ab}-\frac{\bar{R}}{(n-1)}\bar{g}_{ab}\right),\quad 2\leq a,b\leq n,$ (A.4)
@@
A.1. Curvature tensor of cones with vanishing Weyl tensor
@@
where $(\Sigma^{n-1},\bar{g})$ is a closed (n1)(n-1)-dimensional Riemannian manifold. Then, by (A.4), we see that $\mathcal{C}$ has vanishing Weyl curvature $W=0$ if and only if its link $\Sigma$ is a space form. Thus, we immediately have the following
Lemma A.2.
@@
A.2. Curvature decomposition and curvature operator of 4D cones
@@
$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$, (A.5)
@@
$\langle X\wedge Y,V\wedge W\rangle=\langle X,V\rangle\langle Y,W\rangle-\langle X,W\rangle\langle Y,V\rangle.$ (A.6)
Observe that, for the matrices $A$ and $C$ in (A.5), we have
@@
so that $\{e_{1},e_{2},e_{3},e_{4}\}$ forms an orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, by (A.1), we have
@@
Therefore, by (A.7), we obtain
$A_{11}=C_{11}$ $=\tfrac{1}{2}R_{3434}=\tfrac{1}{2r^{2}}(m_{1}-1),$ (A.9)
@@
Similarly, by (A.8), we have
$B_{11}$ $=-\tfrac{1}{2}R_{3434}=-\tfrac{1}{2r^{2}}(m_{1}-1)=-A_{11},$ (A.10)
@@
$R_{11}$ $=0,$ $R_{22}=r^{-2}\left(m_{2}+m_{3}-2\right),$
$R_{33}$ $=r^{-2}\left(m_{1}+m_{3}-2\right),$ $R_{44}=r^{-2}\left(m_{1}+m_{2}-2\right).$
@@
    $(\mathcal{C}^{4},g_{c})$ has half-WPIC if and only if it has WPIC and nonnegative Ricci curvature $Rc\geq 0$, if and only if $\overline{Rc}\geq 2{\bar{g}}$.